# Gradient Inversion Attack on Graph Neural Networks

### v1

@@ 5.2.1 Node Attacker 2-a
The recovery performance of the adjacency matrix for this setting is listed in Table 2. Note that the attacker perfectly recovers the graph structure for the GraphSAGE model. For GCN, even though Proposition 3 states that we can exactly recover the adjacency matrix, we do observe some error in the recovery. This can be potentially attributed to either the limited number of iterations employed in the attack, hindering its convergence, or the optimization procedure encountering a local minimum. Nonetheless, even though not a perfect recovery, the attacker is able to recover the graph structure with high accuracy and precision for GCN.
5.2.2 Node Attacker 2-b
@@
Table 3 shows the results for Node Attacker 2-b. The attacker is capable of obtaining almost perfect recovery of nodal features for the GraphSAGE model. Similarly, for the GCN model, the RNMSE values, although slightly higher, remain sufficiently low, which indicates data leakage. Note that the small error can again be attributed to the potential algorithmic issues such as local optimal as discussed before.
5.2.3 Node Attacker 2-c
@@
In this case, the attacker doesn’t have any prior information about the graph data and tries to recover both the nodal features and the graph structure. The results are shown in Table 4. For the GraphSAGE model, it can be observed that the attacker can recover both the features and the adjacency matrix with high accuracy, in both Facebook and GitHub datasets. Again, such results are consistent with our theoretical results in Proposition 5. Surprisingly, for the GCN framework, the attacker is still able to recover $\mathbf{A}$ with high AUC and AP scores. Note that, no theoretical guarantee was provided in Section 4 for node attacker 2-c for GCN framework. Such empirical results may be attributed to the introduced feature smoothness and sparsity regularizers in equation 14. To further investigate the effect of regularizers, the performance of the attack on the model was evaluated without the regularizers. These results are also presented in Table 4. It can be observed that there is a clear decline in the performance of recovering the adjacency matrix if the regularizers are not introduced, especially in Average Precision (AP). It suggests that the introduced regularizers can effectively improve the performance of the attack.
Table 5: Recovery of $\mathbf{A}$ with Graph Attacker-a.
@@
Table 6: Recovery of $\mathbf{X}$ (RNMSE) Graph Attacker b.
Facebook GitHub
@@ 5.3 Graph Attacker
In this subsection, the performance of different graph attackers are evaluated for attacking nodal features as well as the adjacency matrix of the private subgraph.
5.3.1 Graph Attacker-a
@@
Table 5 lists the recovery performance of the graph structure. It can be observed that the attacker can perfectly recover the adjacency matrix from the gradients of a GraphSAGE model in both datasets. Such results are also consistent with the results of Proposition 6. On the other hand, in the case of GCN, despite the absence of theoretical guarantees, the attacker demonstrates high accuracy, precision, and AUC values in identifying the edges. Such results can again be attributed to the feature smoothness and sparsity regularize which help leverage usual properties for social networks. To assess the impact of the regularizer, we again conduct the attack with and without the regularizer as shown in equations equation 12 and equation 14. The resulting recovered adjacency matrices are visualized through heatmaps as depicted in Figures 3(b) and 3(c). Figure 3(a) illustrates the true graph structure of the private data under attack for comparison. The difference in recovery quality between the two scenarios is evident. When the regularizer is not used, the attack accurately identifies a substantial number of edges. However, this also leads to the erroneous identification of many non-edges as edges, resulting in a lower AP score as also shown in the Figure 3.
@@
Figure 3: The heatmaps of the recovered adjacency matrix with Graph Attacker-a from the gradient of GCN models. (a) actual graph. (b) recovered graph structure with regularizers (AUC=$0.981$, AP=$1.0$). (c) recovered graph structure without regularizers (AUC=$0.977$, AP=$0.86$).
5.3.2 Graph Attacker-b
@@
In this setting, the attacker tries to recover the nodal feature matrix given the gradients and the underlying graph structure. The results are listed in Table 6. It can be observed that while the attacker is unable to recover nodal features from a GCN model, it can recover nodel features from a GraphSAGE model with small errors
The results also corroborate that this attacker is different from Node Attacker 2-b where the gradients of the loss for each node are leaked to the attacker. Unlike this scenario, Node Attacker 2-b can successfully recover the nodal features.
5.3.3 Graph Attacker-c
The results for Graph Attacker-c are listed in Table 7. Similar to the results of Node Attacker 2-c, the adjacency matrix can be accurately recovered in this setting for GraphSAGE. Interestingly, in contrast to the previous results, the attack performs better on the GCN than the GraphSAGE framework. Also, in this setting, the attack fails to reconstruct the nodal features for both GraphSAGE and GCN frameworks.
@@ 5.3.3 Graph Attacker-c
Table 7: Recovery of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (AUC and AP) with Graph Attacker-c.
@@
Table 8: Target node feature recovery $\mathbf{x}_{v}$ (RNMSE ×$\times 10^{-2}$) for Node Attacker 1 in a batched setting.
@@
Table 9: Recovery performance of $\mathbf{X}$ (RNMSE ×$\times 10^{-2}$) for Graph Attacker-b for batched graph dataset.
@@ 82 $\pm$ 27
5.4 Attacking batched data
We also extended the experiments for node and graph attackers to the minibatch setting. Specifically, we only evaluate GraphSAGE for both tasks since it provides the best recovery in the stochastic (single data sample) setting as shown in the previous experiments. Although effective attack techniques Wen et al. (2022) have been proposed to recover image data from large batches, here we tested the performance of the attackers by simply optimizing with respect to the dummy input batch. Hence, the performance may improve if attack techniques for batch data are also incorporated. Due to possible permutation ambiguity in a batch setting, the Hungarian algorithm is used to find the best match with the actual input data. The evaluation metric for the algorithm is the RNMSE between all pairs of the dummy batch and the actual batch. These results are averaged over $10$ independent runs for each batch size. Experiments are conducted for Node Attacker 1 and Graph Attacker-b.
Node Attacker 1
Table 8 shows the performance of recovering the target node features for batches of size $B=5,20$, and $50$. It can be observed that the best results are obtained for the Synthetic Dataset with accurate recovery across all batch sizes. When the batch size is $B=5$, the attack provides accurate recovery across all the $3$ datasets.
Graph Attacker-b
In this scenario, the attack attempts to recover $\mathbf{X}$ for the whole batch with $\mathbf{A}$ being known. To simplify the matching after running the attack we only consider graphs with the same number of nodes. This is done by generating Erdos-Renyi (ER) graphs with the number of nodes $n=50$. For a ER graph with $n$ nodes and edge probability $p$, the average degree of a node is $d=(n-1)p$. We conduct experiments by varying $d$ with different batch sizes. The nodal features are generated by sampling from the standard normal distribution i.e. $\mathcal{N}(0,1)$. Table 9 lists the performance of recovering $\mathbf{X}$ with varying batch sizes and degrees. As expected, the quality of recovery declines with increasing batch size, with the best results being achieved at $B=5$. Also, the performance remains roughly the same across all degrees.
5.5 Sensitivity Analysis for Hyperparameters
@@
Figure 4: Variation in RNMSE of the recovered target features vs. the number of hidden channels for a GCN framework on the Github dataset.
Network Width
@@
All previous experiments were conducted with the number of hidden channels set to 100. In the setting of Node Attacker 1, the attack could successfully recover the target nodal features for a GraphSAGE framework with high accuracy. However, for a GCN framework, the attack only managed to recover the aggregated nodal features. Consequently, in this section, we explore the impact of varying the number of hidden channels in the 2-layer GCN framework on the quality of recovery. Figure 4 illustrates the results of target node feature recovery for the Github Dataset. Notably, as the number of hidden channels increases, there is a clear decrease in the RNMSE values, suggesting that wider networks are more susceptible to the attack.
Figure 5: Variation of AUC and AP vs. $\alpha$ and $\beta$ for Graph Attacker-a.
Regularization Hyperparameters
@@
To investigate the influence of $\alpha$ and $\beta$ in equation 14, we examine the effects on the recovery performance of Graph Attacker-a for the GCN framework. More specifically, we maintain one hyperparameter at a fixed value while altering the other, and depict the AUC and AP values. Figure 5(a) illustrates the AUC and AP values for different $\alpha$ values in a log scale with base $10$. The values on the x-axis are also in the log scale with base $10$. In Figure 5(a), $\alpha$ is varied while keeping $\beta=10^{-7}$. Similarly, in Figure 5(b), where $\beta$ is varied, $\alpha$ is fixed at $10^{-9}$. It is evident from the figures that higher values of hyperparameters severely degrade the performance of the attack whereas extremely low values give suboptimal results. This clearly demonstrates the effectiveness of the feature smoothness and sparsity regularizers.
5.6 Numerical Tests on Molecular Datasets
All attack settings considered in previous experiments are defined within the context of social networks. In order to assess the efficacy of the attack on alternative datasets, such as molecular datasets which are commonly employed in graph classification tasks, we extend our analysis to include three datasets from the TUDataset Morris et al. (2020). The experimental setting, evaluation metrics and results are given in detail in Appendix F. The average number of nodes in the graphs of these datasets is significantly smaller than social networks (see Table 11). However, unlike social networks, these graphs might not necessarily exhibit properties like feature smoothness and sparsity. Hence, for all the attacks we use the objective function as defined in equation 12. To improve the attack for Graph Attacker-c where we have no prior knowledge of the graph we utilize various initialization and thresholding strategies. Specifically, since the optimization process given in the algorithms can be greatly improved by utilizing different starting points, we initialize the dummy node feature and adjacency matrices with different values. In the case of thresholding, instead of sampling the edges given the probabilistic adjacency matrix $\hat{\mathbf{A}}^{P+1}$, we instead use various threshold values to sample edges. For instance, if the threshold is set to $0.5$ we only sample edges from the probabilistic adjacency matrix $\hat{\mathbf{A}}^{P+1}$ with values greater than $0.5$. The results are listed in Table 15, which demonstrate that even in this case, without any consideration about sparsity or smoothness, while simply utilizing different initializations and thresholds, the attacker can still recover the nodal features as well as the underlying graph structure with high accuracy. The results in the other attack settings are again consistent with the theoretical analysis. For example, for Graph Attacker-a, the underlying graph structure recovery is only possible for one of the datasets, see Table 5(d). Note that the nodal feature matrix $\mathbf{X}$ was not full-row rank for the MUTAG and COIL-RAG datasets. Such results were consistent with Proposition 6, which states that the underlying graph structure can be recovered only when $\mathbf{X}$ is full-row rank.
6 Conclusion & Future Work
In this work, we presented the first attempt to study the gradient inversion attack for Graph Neural Networks(GNNs) in a federated graph learning setting and explored what information can be leaked about the graph data from the gradient. We studied both the node classification and the graph classification tasks, along with two widely used GNN structures (GraphSAGE and GCN). Through both theoretical and empirical results, we highlighted and analyzed the vulnerabilities of different GNN frameworks across a wide variety of gradient inversion attack settings.
Future work
In the future, it would be interesting to study the extension of the attack mechanisms to other GNN frameworks, such as GIN (Graph Isomorphism Networks) Xu et al. (2019) and GAT (Graph Attention Networks)Veličković et al. (2018). Experiments and theoretical analysis regarding the vulnerability of other GNN frameworks hence is a promising direction. For example, are certain GNN frameworks more robust to gradient inversion attacks than others? Also, are there any other special properties that can be utilized for graph datasets other than social networks? For instance, it is worth exploiting the fact that the degrees of nodes in molecular datasets exhibit an upper limit. Finally, recent studies such as Wang et al. (2023) have provided theoretical bounds on the recovery quality of minibatch data. Hence, an intriguing question would be whether similar guarantees can be established for graph data.
Furthermore, it is crucial to have appropriate defense mechanisms in place to prevent data leakage. It is commonly believed that employing larger batch sizes could be an effective defense against the gradient inversion attack, but our experiments show that even in batches of size $50$, the private data can be recovered with high accuracy. Therefore, solely relying on larger batch sizes is insufficient as a defense strategy against gradient inversion attacks. Recent studies such as Wen et al. (2022); Wang et al. (2023) have provided provable guarantees regarding the quality of recovered private data from training batches. Based on the various defense strategies introduced in Gupta et al. (2022), some potential defense mechanisms can remain effective in graph setting: e.g., noisy gradients Abadi et al. (2016), gradient pruning Zhu et al. (2019), encoding inputs Huang et al. (2020). However, more specific defensive mechanisms from graph perspectives remain undeveloped.
Appendix A Dataset
Table 10: Dataset statistics (Node Classification).
Dataset #Nodes #Edges #Labels
GitHub 37,300 578,006 2
Facebook 22,470 342,004 4
Appendix B Extracting ground-truth labels
$=g_{i}\cdot\lambda_{i}\cdot\mathbf{h}_{v}^{\text{agg}}.$ (25)
Appendix C Proof of Proposition 1 for GraphSAGE
$\mathbf{h}_{v}$ $=\sigma(\tilde{\mathbf{h}}_{v}),$ (27)
Appendix D Proof of Proposition 2
$\mathbf{h}_{v}$ $=\sigma(\tilde{\mathbf{h}}_{v}),$ (31)
The target nodal features can henceforth be recovered as $\mathbf{x}_{v}=\nabla_{(\mathbf{W}_{2})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
Appendix E Proof of Proposition 6
Proof.
$\mathbf{H}$ $=\sigma(\tilde{\mathbf{H}}),$ (35)
Appendix F Additional Graph Classification results on TUDatasets

### v2

@@ 6.2 Node Attacker 2
The reconstruction performance of the adjacency matrix for this setting is listed in Table 4. Note that the attacker perfectly reconstructs the graph structure for the GraphSAGE model. For GCN, even though Proposition 3 states that we can exactly reconstruct the adjacency matrix, we do observe some error in the reconstruction. This can be potentially attributed to either the limited number of iterations employed in the attack, hindering its convergence, or the optimization procedure encountering a local minimum. Nonetheless, even though not a perfect reconstruction, the attacker can reconstruct the graph structure with high accuracy and precision for GCN. GLG consistently outperforms DLG for both GraphSAGE and GCN, demonstrating its overall superiority in reconstructing the adjacency matrix.
Node Attacker 2-N.
@@
Table 5 shows the results for Node Attacker 2-N. The attacker is capable of obtaining almost perfect reconstruction of nodal features for the GraphSAGE model. Similarly, for the GCN model, the RNMSE values, although slightly higher, remain sufficiently low, which indicates data leakage. Note that the small error can again be attributed to the potential algorithmic issues such as local optimal as discussed before. Compared to DLG, GLG consistently performs better in reconstructing nodal features for GCN. While DLG achieves slightly better results on the GitHub dataset with GraphSAGE, like Node Attacker 1, this isolated exception does not detract from the overall trend of GLG’s reliable performance across settings.
Node Attacker 2-GN.
@@
In this experimental setting, the attacker possesses no prior information regarding the graph data and endeavors to reconstruct both the nodal features and the graph structure. Table 6 summarizes the corresponding results. For the GraphSAGE model, it is evident that the adversary can reconstruct both the features and the adjacency matrix with high accuracy on both the Facebook and GitHub datasets. These outcomes are consistent with our theoretical findings presented in Proposition 5. Notably, for the GCN framework, the adversary is still able to reconstruct $\mathbf{A}$ with high AUC and AP scores, despite the absence of a theoretical guarantee for node attacker 2-GN in the GCN context (as noted in Section 5). This empirical observation may be attributed to the integration of the feature smoothness and sparsity regularizers defined in Equation equation 14.
To further evaluate the impact of the regularizers, we assessed the performance of the attack without their inclusion. The corresponding results, also presented in Table 6, indicate a marked decline in the reconstruction performance of the adjacency matrix, particularly in terms of Average Precision (AP), when the regularizers are omitted. This suggests that the inclusion of these regularizers effectively enhances the performance of the attack.
Furthermore, we compare the performance of DLG with our proposed method, GLG. Our empirical results demonstrate that DLG is capable of reconstructing both the nodal features and the graph structure only within the GraphSAGE framework, in accordance with Proposition 5. In contrast, for the GCN model on both the Facebook and GitHub datasets, GLG consistently outperforms DLG, particularly with respect to the reconstruction of the adjacency matrix as indicated by the AUC and AP metrics. These findings underscore the broader applicability and superior performance of GLG, which is effective even in scenarios lacking formal theoretical guarantees.
Table 7: Reconstruction of $\mathbf{A}$ with Graph Attacker-G.
@@
DLG
GraphSAGE 0.50$\pm$0.01 0.50$\pm$0.03 0.13$\pm$0.05 0.50$\pm$0.04 0.51$\pm$0.02 0.15$\pm$0.04
GCN 0.51$\pm$0.01 0.51$\pm$0.02 0.17$\pm$0.07 0.53$\pm$0.01 0.53$\pm$0.02 0.15$\pm$0.04
Table 8: Reconstruction of $\mathbf{X}$ (RNMSE) with Graph Attacker-N.
Framework Facebook GitHub
GLG (Ours)
@@ 6.3 Graph Attacker
In this subsection, the performance of different graph attackers is evaluated for attacking nodal features as well as the adjacency matrix of the private subgraph.
Graph Attacker-G.
@@
Table 7 lists the reconstruction performance of the graph structure. It can be observed that the attacker can perfectly recover the adjacency matrix from the gradients of a GraphSAGE model in both datasets. Such results are also consistent with the results of Proposition 6. On the other hand, in the case of GCN, despite the absence of theoretical guarantees, the attacker demonstrates high accuracy, precision, and AUC values in identifying the edges. Such results can again be attributed to the feature smoothness and sparsity regularizer which help leverage usual properties for social networks. Unlike the Node attacker setting, DLG performs poorly in reconstructing the graph structure under the graph attacker setting, highlighting its limitations in directly handling graph-structured data. To assess the impact of the regularizer, we again conduct the attack with and without the regularizer as shown in equations equation 12 and equation 14. The resulting recovered adjacency matrices are visualized through heatmaps as depicted in Figures 2(b) and 2(c). Figure 2(a) illustrates the true graph structure of the private data under attack for comparison. The difference in reconstruction quality between the two scenarios is evident. When the regularizer is not used, the attack accurately identifies a substantial number of edges. However, this also leads to the erroneous identification of many non-edges as edges, resulting in a lower AP score as also shown in the Figure 2.
@@
Figure 2: The heatmaps of the reconstructed adjacency matrix with Graph Attacker-G from the gradient of GCN models. (a) actual graph. (b) reconstructed graph structure with regularizers (AUC=0.9810.9810.981, AP=1.01.01.0). (c) reconstructed graph structure without regularizers (AUC=0.9770.9770.977, AP=0.860.860.86).
Graph Attacker-N.
@@
In this setting, the attacker tries to reconstruct the nodal feature matrix given the gradients and the underlying graph structure. The results are listed in Table 8. It can be observed that while the attacker is unable to reconstruct nodal features from a GCN model, it can recover nodal features from a GraphSAGE model with small errors. The results also corroborate that this attacker is different from Node Attacker 2-N where the gradients of the loss for each node are leveraged to the attacker. Unlike this scenario, Node Attacker 2-N can successfully reconstruct the nodal features. While GLG performs better on GraphSAGE than on GCN, it still consistently outperforms DLG on both architectures, suggesting that DLG struggles to directly handle graph-structured data well due to interdependencies between nodal features and graph topology.
Graph Attacker-GN.
Table 9 reports the performance of the Graph Attacker-GN in reconstructing the graph structure and nodal features. An analysis of the results for our proposed method, GLG, reveals that the reconstructed adjacency matrix is obtained with high accuracy for the GraphSAGE framework. Interestingly, in this experimental setting, the attack on the GCN framework yields even better performance in recovering the graph structure, as compared to GraphSAGE. It is worth noting, however, that in both frameworks the reconstruction of the nodal features remains unsatisfactory, highlighting an intrinsic difficulty in recovering node attributes under this attack paradigm.
A detailed comparison between GLG and the baseline method, DLG, further underscores the superiority of GLG. In the GraphSAGE setting, while DLG is able to some success reconstruction of the graph structure (for Github), its performance is still far inferior to that of GLG in terms of key metrics such as RNMSE, AUC, and AP. Moreover, even under the GCN framework, GLG exceeds the performance of DLG, despite the absence of theoretical guarantees for nodal feature reconstruction. These empirical findings demonstrate that GLG is more robust and effective in adversarial graph reconstruction, providing enhanced accuracy in reconstructing the adjacency matrix even in challenging scenarios.
Overall, the results clearly establish that GLG is a more reliable and superior approach compared to DLG, as it consistently achieves higher reconstruction accuracy across different graph models and datasets.
@@
Table 9: Reconstruction of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (AUC and AP) with Graph Attacker-GN for two different methodologies on Facebook and GitHub datasets.
@@
DLG
GraphSAGE 1.39$\pm$0.01 0.50$\pm$0.02 0.20$\pm$0.08 0.85$\pm$0.73 0.71$\pm$0.25 0.48$\pm$ 0.46
GCN 1.31$\pm$0.05 0.58$\pm$0.16 0.24$\pm$0.14 0.78$\pm$0.06 0.88$\pm$0.05 0.70$\pm$0.08
7 Conclusion & Future Work
In this work, we study the gradient inversion attack for Graph Neural Networks (GNNs) in a federated graph learning setting and explore what information can be leaked about the graph data from the gradient. We studied both the node classification and the graph classification tasks, along with two widely used GNN structures, GraphSAGE and GCN. Through both theoretical and empirical results, we highlighted and analyzed the vulnerabilities of different GNN frameworks across a wide variety of gradient inversion attack settings. Our proposed method, GLG, attains superior accuracy in reconstructing both nodal features and graph structure from gradients.
Future work.
In the future, it would be interesting to study the extension of the attack mechanisms to other GNN frameworks, such as GIN (Graph Isomorphism Networks) (Xu et al., 2019) and GAT (Graph Attention Networks) (Veličković et al., 2018). Experiments and theoretical analysis regarding the vulnerability of other GNN frameworks hence is a promising direction. For example, are certain GNN frameworks more robust to gradient inversion attacks than others? Also, are there any other special properties that can be utilized for graph datasets other than social networks? For instance, it is worth exploiting the fact that the degrees of nodes in molecular datasets exhibit an upper limit. Finally, recent studies such as Wang et al. (2023) have provided theoretical bounds on the reconstruction quality of minibatch data. Hence, an intriguing question would be whether similar guarantees can be established for graph data.
Furthermore, it is crucial to have appropriate defense mechanisms in place to prevent data leakage. It is commonly believed that employing larger batch sizes could be an effective defense against the gradient inversion attack, but our experiments show that even in batches of size 505050, the private data can be reconstructed with high accuracy. Therefore, solely relying on larger batch sizes is insufficient as a defense strategy against gradient inversion attacks. Recent studies such as Wen et al. (2022); Wang et al. (2023) have provided provable guarantees regarding the quality of reconstructed private data from training batches. Based on the various defense strategies introduced in Gupta et al. (2022), some potential defense mechanisms can remain effective in graph setting: e.g., noisy gradients (Abadi et al., 2016), gradient pruning (Zhu et al., 2019), encoding inputs (Huang et al., 2020). However, more specific defensive mechanisms from graph perspectives remain undeveloped.
Acknowledgments
Work in the paper is supported by NSF ECCS 2412484, ECCS 2207457, NSF GEO CI 2425748, and NSF SaTC Frontiers FG22920.
Appendix A Attack Algorithms
Algorithm 1 GLG (Node Attacker).
1:Input: 222-layer GNN model $F(\mathbf{x}_{v},\{\mathbf{x}_{v_{j}}\}_{j\in\mathcal{N}_{v}},\mathbf{W})$
[…]
3:Output: private data $\mathbf{x}_{v}$
6:for $p\leftarrow 1\textrm{ to }P$ do
7:///// /Compute dummy gradients:
[…]
9: Compute $\mathbb{D}^{p}$ as in equation 12
[…]
11:for $j\in\mathcal{N}_{v}$ do
13:end for
14:end for
15:return $\hat{\mathbf{x}}=\hat{\mathbf{x}}_{v}^{P+1}$
Algorithm 2 GLG (Node Attacker 2)
2:Output: private training data $\mathbf{X},\mathbf{A}$
4:for $p\leftarrow 1\textrm{ to }P$ do
5:///// /Compute dummy gradients:
[…]
7: Compute $\hat{\mathbb{D}}^{p}$ as given in equation 14
[…]
10:end for
Algorithm 3 Graph Attacker
2:Output: private training data $\mathbf{X},\mathbf{A}$
4:for $p\leftarrow 1\textrm{ to }P$ do
5:///// /Compute dummy gradients:
[…]
7: Compute $D^{p}$ as given in equation 12
[…]
10:end for
Appendix B Extracting ground-truth labels
$=g_{i}\cdot\lambda_{i}\cdot\mathbf{h}_{v}^{\text{agg}}.$ (25)
Appendix C Proof of Proposition 1 for GraphSAGE
$\mathbf{h}_{v}$ $=\sigma(\tilde{\mathbf{h}}_{v}),$ (27)
Appendix D Proof of Proposition 2
$\mathbf{h}_{v}$ $=\sigma(\tilde{\mathbf{h}}_{v}),$ (31)
The target nodal features can henceforth be reconstructed as $\mathbf{x}_{v}=\nabla_{(\mathbf{W}_{2})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
Appendix E Proof of Proposition 6
Proof.
$\mathbf{H}$ $=\sigma(\tilde{\mathbf{H}}),$ (35)
Appendix F Supplementary Experimental Results and Analysis
Table 10: Dataset statistics for Node Classification task.
Dataset #Nodes #Edges #Labels
GitHub 37,300 578,006 2
Facebook 22,470 342,004 4
F.1 Node Attacker 1: One-hop Neighborhood Nodal Feature Reconstruction
To evaluate the performance of the node attacker in reconstructing the nodal features of one-hop neighbors, a target node is selected randomly from the graph. A dummy graph is initialized as a depth-222 tree with $d_{tree}$ set to the actual number of one-hop neighbors of the target node. Since there’s permutation ambiguity in the reconstructed nodal features, the Hungarian Algorithm is used to find the best match with the original neighboring node features (Kuhn, 2010). Figure 3 shows the one-hop neighbor reconstruction error (RNMSE) using GraphSAGE. The results are averaged over 101010 randomly selected nodes. For some nodes, the nodal features of neighbors were reconstructed accurately. For example, nodes 0 and 888 in Figure 3(a) for Facebook. Further, although the mean RNMSE of the reconstructed neighbors is high, the minimum RNMSE of the best-reconstructed neighbors can be low (node 999 in Figure 3(a)), indicating private data leakage.
Figure 3: One-hop neighbor reconstruction (RNMSE) for Node Attacker-1 with GraphSAGE.
F.2 Attacking batched data
Table 11: Target node feature reconstruction $\mathbf{x}_{v}$ (RNMSE ×$\times 10^{-2}$) for Node Attacker 1 in a batched setting.
@@ F.2 Attacking batched data
Table 12: Reconstruction performance of $\mathbf{X}$ (RNMSE ×$\times 10^{-2}$) for Graph Attacker-N for batched graph dataset.
@@ 82 $\pm$ 27
We also extended the experiments for node and graph attackers to the minibatch setting. Specifically, we only evaluate GraphSAGE for both tasks since it provides the best reconstruction in the stochastic (single data sample) setting as shown in the previous experiments. Although effective attack techniques (Wen et al., 2022) have been proposed to reconstruct image data from large batches, here we tested the performance of the attackers by simply optimizing with respect to the dummy input batch. Hence, the performance may improve if attack techniques for batch data are also incorporated. Due to possible permutation ambiguity in a batch setting, the Hungarian algorithm is used to find the best match with the actual input data. The evaluation metric for the algorithm is the RNMSE between all pairs of the dummy batch and the actual batch. These results are averaged over 101010 independent runs for each batch size. Experiments are conducted for Node Attacker 1 and Graph Attacker-N.
Node Attacker 1.
Table 11 shows the performance of reconstructing the target node features for batches of size $B=5,20$, and 505050. It can be observed that the best results are obtained for the Synthetic Dataset with accurate reconstruction across all batch sizes. When the batch size is $B=5$, the attack provides accurate reconstruction across all the 333 datasets.
Graph Attacker-N.
In this scenario, the attack attempts to reconstruct $\mathbf{X}$ for the whole batch with $\mathbf{A}$ being known. To simplify the matching after running the attack we only consider graphs with the same number of nodes. This is done by generating Erdos-Renyi (ER) graphs with the number of nodes $n=50$. For a ER graph with $n$ nodes and edge probability $p$, the average degree of a node is $d=(n-1)p$. We conduct experiments by varying $d$ with different batch sizes. The nodal features are generated by sampling from the standard normal distribution i.e. $\mathcal{N}(0,1)$. Table 12 lists the performance of reconstructing $\mathbf{X}$ with varying batch sizes and degrees. As expected, the quality of reconstruction declines with increasing batch size, with the best results being achieved at $B=5$. Also, the performance remains roughly the same across all degrees.
F.3 Sensitivity Analysis for Hyperparameters
@@
Figure 4: Variation in RNMSE of the reconstructed target features vs. the number of hidden channels for a GCN framework on the Github dataset.
Network Width.
@@
All previous experiments were conducted with the number of hidden channels set to 100. In the setting of Node Attacker 1, the attack could successfully reconstruct the target nodal features for a GraphSAGE framework with high accuracy. However, for a GCN framework, the attack only managed to reconstruct the aggregated nodal features. Consequently, in this section, we explore the impact of varying the number of hidden channels in the 2-layer GCN framework on the quality of reconstruction. Figure 4 illustrates the results of target node feature reconstruction for the Github Dataset. Notably, as the number of hidden channels increases, there is a clear decrease in the RNMSE values, suggesting that wider networks are more susceptible to the attack.
Figure 5: Variation of AUC and AP vs. $\alpha$ and $\beta$ for Graph Attacker-G.
Regularization Hyperparameters.
@@
To investigate the influence of $\alpha$ and $\beta$ in equation 14, we examine the effects on the reconstruction performance of Graph Attacker-G for the GCN framework. More specifically, we maintain one hyperparameter at a fixed value while altering the other, and depict the AUC and AP values. Figure 5(a) illustrates the AUC and AP values for different $\alpha$ values in a log scale with base 101010. The values on the x-axis are also in the log scale with base 101010. In Figure 5(a), $\alpha$ is varied while keeping $\beta=10^{-7}$. Similarly, in Figure 5(b), where $\beta$ is varied, $\alpha$ is fixed at $10^{-9}$. It is evident from the figures that higher values of hyperparameters severely degrade the performance of the attack whereas extremely low values give suboptimal results. This clearly demonstrates the effectiveness of the feature smoothness and sparsity regularizers.
Appendix G Additional Graph Classification results on TUDatasets
All attack settings considered in previous experiments in Section 6 are defined within the context of social networks. In order to assess the efficacy of the attack on alternative datasets, such as molecular datasets which are commonly employed in graph classification tasks, we extend our analysis to the TUDataset (Morris et al., 2020).